# The Impact of Unstated Norms in Bias Analysis of Language Models

### v4

@@ 5 Discussion
English pretraining data for LLMs is dominated by text drawn from areas where the racial majority is White Bender et al. (2021); Navigli et al. (2023). Several studies have confirmed that markedness is widespread in internet data, with White race and male gender constituting the unmarked defaults Wolfe and Caliskan (2022c); Bailey et al. (2022). Furthermore, it has been shown that models, and LLMs in particular, trained on web data reflect these markedness characteristics Bender et al. (2021); Wolfe and Caliskan (2022a, b). On the other hand, in templates commonly used for bias quantification, race is explicitly mentioned to establish group membership. As such, template-based text that explicitly establishes that the subject is “White” essentially constitute out-of-domain examples Blodgett et al. (2021); Dressler (1985). Such a mismatch likely influences model predictions. Thus, model errors or disparities with respect to unmarked groups are unlikely to be natural biases, but instead reflect a deficiency in probe design. The results above demonstrate that these probes can produce misleading measurements that need to be considered more deeply.
While these findings are unexpected, some evidence of this issue exists in other studies but has not been systematically investigated. In Brown et al. (2020), sentiment scores for GPT-3 are generally lower when explicitly linked to White race compared with others, including Black race. The authors hypothesize that the sentiment associated with an “unspecified” race, more accurately represents the sentiment associated with White race. It has also been observed that Gopher-280B tends to classify text examples from “Black,” “White,” and “Muslim” groups as toxic at a higher rate than others Rae et al. (2021). Finally, in Bai et al. (2022), the average sentiment of generations from base models and models after alignment are compared. While alignment generally improves sentiment across all groups, “Black” and “White” generations have the lowest sentiment in all cases.
The results in this work, coupled with previous evidence, highlight the need to consider the impact that the use of bias probes that rely on marked text has on the measurement of bias. In this case, such probes produce largely misleading results. Ideally, artificial injection of demographic information would not be required. For example, the studies of Seyyed-Kalantari et al. (2020) and Sap et al. (2019) establish group membership through meta-data, self-identification, or classification techniques rather than explicitly in text. These methods avoid the out-of-domain nature of template-based examples of the kind studied here and do not see the unnatural patterns we observed.
Alternatively, the inclusion of datasets explicitly correcting for markedness in the pretraining of LLMs could help better align template-based text. Table 4 provides a small indication that the newest LLMs, trained on increasingly large and multilingual datasets, show improvements in the overall size of the measured gaps. Both Llama-3-8B and Mistral-7B have the smallest difference between the most positive and negative gaps for Negative-Sentiment FPR, averaged over the three datasets. Llama-3-8B also produces the lowest average value for this difference for Positive-Sentiment FPR. Given that the gaps for the White group often represent one of the extreme gaps in the experiments, this suggests that newer models might also be less affected by markedness.
In this paper, several unexpected bias measurements with respect to race were presented. Rather than representing a true social bias in LLMs, it is conjectured that the observed outliers are due to a misalignment of the template-based bias probes and the underlying LLM pretraining data due to markedness. These results highlight an important weakness in the commonly applied approach to counterfactual bias quantification. In future work, we aim to design experiments to further confirm the misalignment conjecture. Additionally, we intend to investigate whether multimodal models can overcome the issues highlighted here.

### v5

@@ Appendix C Fine-tuning and Prompting Accuracy
Appendix D Gap Differences Across Models
For each of the fine-tuned models, across the different datasets, an FPR-gap span is calculated. For a given type of FPR, Negative- or Positive-Sentiment gap spans are computed as the largest difference between any two mean FPR gaps for the groups. This quantifies how large the particular FPR disparities for a given model and dataset are between groups. The larger this span, the greater the difference in Negative- or Positive-Sentiment FPR between groups and the less invariant the model is to overall group substitution. Table 4 displays the FPR gap spans for each model, averaged over the three datasets. In computing the spans, the gap for the White group is part of the span extremes 58% of the time for Negative-Sentiment FPR and 100% of the time for Positive-Sentiment FPR. That is, the gap computed for the White group often constitutes one of the largest gap magnitudes.
From the table, it is clear that the RoBERTa and Llama-2 models have consistently large spans for both types of FPR gap. On the other hand, Llama-3-8B, the most recent model studied with fine-tuning, has the smallest average gap spans in both categories. Another recent model, Mistral-7B, demonstrates a small average Negative-Sentiment FPR gap span, suggesting that more recent LLMs may be slightly less affected by issues with the template-based probes. It is interesting to note that the distribution of spans for Positive-Sentiment FPR gaps are more uniformly distributed between models than the Negative-Sentiment counterpart.
Table 5: Statistics for the three template-based probing datasets, Amazon, Regard, and NS-Prompts. The label counts and distributions are reported for each racial group across datasets. The label distributions are constant between groups, and thus reported below the label totals. Note that all examples from the NS-Prompts dataset have a neutral label by construction.
Label Counts
Dataset Race Negative Neutral Positive Total Fraction
Amazon African American $200$ $200$ $200$ $600$ $0.182$
American Indian $300$ $300$ $300$ $900$ $0.273$
Asian $100$ $100$ $100$ $300$ $0.091$
Hispanic $200$ $200$ $200$ $600$ $0.182$
Pacific Islander $200$ $200$ $200$ $600$ $0.182$
White $100$ $100$ $100$ $300$ $0.091$
Total $1100$ $1100$ $1100$ $3300$
Distribution $0.333$ $0.333$ $0.333$
Regard African American $440$ $220$ $350$ $1010$ $0.182$
American Indian $660$ $330$ $525$ $1515$ $0.273$
Asian $220$ $110$ $175$ $505$ $0.091$
Hispanic $440$ $220$ $350$ $1010$ $0.182$
Pacific Islander $440$ $220$ $440$ $1010$ $0.182$
White $220$ $110$ $175$ $505$ $0.091$
Total $2420$ $1210$ $1925$ $5555$
Distribution $0.437$ $0.218$ $0.347$
NS-Prompts African American 0 $8880$ 0 $8880$ $0.182$
American Indian 0 $13320$ 0 $13320$ $0.273$
Asian 0 $4440$ 0 $4440$ $0.091$
Hispanic 0 $8880$ 0 $8880$ $0.182$
Pacific Islander 0 $8880$ 0 $8880$ $0.182$
White 0 $4440$ 0 $4440$ $0.091$
Total 0 $48840$ 0 $48840$
Distribution $0.0$ $1.0$ $0.0$
Appendix E Template-Based Dataset Statistics
In this appendix, some statistics and distribution information about the template-based probing datasets described in Section 3.1 are presented. Table 5 summarizes the label counts across datasets and broken down by racial groups. Note that the label distributions are constant between groups. For example, the label distribution of $(0.333,0.333,0.333)$ holds for each protected group as well as the aggregated label distribution. Further, it is important to recall that the models are not trained on these datasets and only perform classification inference to produce the results of Section 4. For the Amazon dataset, the labels are balanced over the entire dataset and within each individual group. The distribution for the Regard dataset is also fairly balanced with somewhat fewer neutral examples. Finally, as discussed in Section 3.1.2, all examples for the NS-Prompts dataset have a neutral label.
Appendix F Supporting Results for Unmarked Groups in Other Sensitive Attributes
This work primarily focuses on unexpected measurements with respect to the sensitive attribute of race. However, the conjecture that markedness causes, or at least contributes to, the observed irregularities when using template-based probes would, in theory, generalize beyond racial demographics. For example, as discussed in Section 5.1, several previous studies have noted that gender-based markedness is also prevalent in web data and influences model behaviors. In such settings, male gender represents the unmarked default. Furthermore, linguistic markedness extends to sexuality, with heterosexuality comprising the predominant unmarked demographic (Zerubavel, 2018).
With this in mind, two additional experiments are conducted using the prompting approaches described in Section 3.4 and the Amazon dataset (Czarnowska et al., 2021). The Amazon dataset includes templates for sensitive attributes beyond race, including sexuality and gender, with the same structure described in Section 3.1.1. Prompt-based classification is applied to these templates, and the same FPR gaps are measured across various protected groups within the sensitive attributes.
Figure 5: Negative- and Positive-Sentiment FPR gaps for protected group variations within the sensitive attribute of Sexuality as measured by the Amazon dataset. In the legend, model names without a suffix indicate zero-shot prompting. SST5 and SE indicate 9-shot prompts with examples drawn from the SST5 and SemEval datasets, respectively.
The results are shown in Figures 5 and 6. For sexuality, the traditionally unmarked group, heterosexual, reflects a similar FPR-gap pattern to the White group in the main results. That is, positive, and often statistically significant, Negative-Sentiment FPR gaps and negative Positive-Sentiment FPR gaps are found. Moreover, these patterns correlate to traditionally underprivileged groups in homosexual and asexual orientation for the respective FPR-gap types. When considering gender, the templates associated with male gender also produce positive and negative gaps for Negative- and Positive-Sentiment FPR, respectively, though the frequency of statistical significance is slightly reduced.
As noted in Section 5.1, additional experimentation is required to confirm that markedness is a driving factor for the results presented in this paper. Nonetheless, the empirical results in this section reinforce the observation that template-based probes produce imperfect measures of bias in LLMs and that these imperfections appear to affect measurements associated with traditionally unmarked groups, even beyond of the sensitive attribute of race.
Figure 6: Negative- and Positive-Sentiment FPR gaps for protected group variations within the sensitive attribute of Gender as measured by the Amazon dataset. In the legend, model names without a suffix indicate zero-shot prompting. SST5 and SE indicate 9-shot prompts with examples drawn from the SST5 and SemEval datasets, respectively.